OpenStudy (anonymous):

The whiteboard in the SPA has a perimeter of 408 inches. The length is 11 inches less than 4 times the height. How many inches high is the whiteboard?

OpenStudy (whpalmer4):

perimeter (P) = 2*L + 2*H P = 408 L = 4*H - 11 Substitute formula for L in terms of H into formula for P and set equal to known perimeter. Solve for H.

OpenStudy (anonymous):